Good evening ma'am or sir, would you please help me on how to solve these 3 equations :
1). (ax^3)^2(ax^4)^3
2). (ab^2)^4(a^4b^2)^3
3). (5a^2xy^3)(b^3y^4)^3
Thank you very much ma'am or sir for your generosity. :)

Answer by checkley71(8403)

(AX^3)^2=A^2X^6
(AX^4)^3=A^3X^12
(A^2X^6)(A^3X^12)=A^5X^18 ANSWER.